# Homogenized Equations for Isentropic Gas in a Pipe with Periodically-Varying Cross-Section

### v1

@@ 1 Introduction
The main goal of this work is to construct an effective medium that represents the essential characteristics of the original microstructured medium. To accomplish this, we develop a constant-coefficient (homogenized) approximation, which effectively describes the behavior of long-wavelength solutions for a system that describes the scenario of a gas propagating in a narrow pipe with a periodically-varying cross-sectional area. This approximation method draws upon the approach developed by Yong and co-authors [18, 8], and more recently was applied to the shallow water waves over periodically-varying bathymetry [17].
@@ 2.2 Averaging Operators
[…]
@@ 2.3.3 ${\mathcal{O}}(\delta^{1})$
00 0 $=a\underline{\rho}^{(1)}_{t}+\underline{q}^{(1)}_{x}+q^{(2)}_{y},$ (2.10a)
00 0 $=-\frac{a^{-2}a_{y}}{\underline{\rho}^{(0)}}(\underline{q}^{(1)})^{2}+\underline{q}^{(1)}_{t}+aP^{\prime}(\underline{\rho}^{(0)})\left(\underline{\rho}^{(1)}_{x}+\rho^{(2)}_{y}\right).$ (2.10b)
@@
[…]
@@
[…]
@@
$\underline{\rho}_{t}$ $=-\frac{\underline{q}_{x}}{\langle{a}\rangle}+{\mathcal{O}}(\delta^{2}),$ (2.14a)
$\underline{q}_{t}$ $=-\frac{P^{\prime}(\underline{\rho}^{(0)})}{\langle{a^{-1}}\rangle}\underline{\rho}_{x}+{\mathcal{O}}(\delta^{2}).$ (2.14b)
@@ 2.3.4 ${\mathcal{O}}(\delta^{2})$
00 0 $=a\rho^{(2)}_{t}+q^{(2)}_{x}+q^{(3)}_{y},$ (2.16a)
00 0 $=q^{(2)}_{t}+a^{-2}a_{y}\left(-\frac{2q^{(1)}q^{(2)}}{\rho^{(0)}}+\frac{(q^{(1)})^{2}\rho^{(1)}}{(\rho^{(0)})^{2}}\right)+2a^{-1}\left(\frac{q^{(1)}q^{(2)}_{y}+q^{(1)}q^{(1)}_{x}}{\rho^{(0)}}\right)$
@@
[…]
@@
$\underline{\rho}_{t}$ $=-\frac{\underline{q}_{x}}{\langle{a}\rangle}+\delta\left(-\frac{2\langle{a\llbracket{a^{-3}a_{y}}\rrbracket}\rangle}{\langle{a^{-1}}\rangle\underline{\rho}^{(0)}\langle{a}\rangle}\underline{q}\,\underline{\rho}_{x}-\frac{\langle{a\llbracket{a^{-1}}\rrbracket}\rangle}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}\underline{q}_{xx}\right)+{\mathcal{O}}(\delta^{3}),$ (2.19a)
$\underline{q}_{t}$ $=-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}\underline{\rho}_{x}+\delta\Biggr{[}\left(\frac{-2}{\underline{\rho}^{(0)}\langle{a}\rangle}+\frac{2\langle{a^{-3}a_{y}\llbracket{a}\rrbracket}\rangle}{\langle{a^{-1}}\rangle\underline{\rho}^{(0)}\langle{a}\rangle}\right)\underline{q}\,\underline{q}_{x}-\frac{P^{\prime\prime}}{\langle{a^{-1}}\rangle}\underline{\rho}\,\underline{\rho}_{x}+\frac{\langle{a^{-1}\llbracket{a}\rrbracket}\rangle P^{\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}\underline{\rho}_{xx}\Biggl{]}+{\mathcal{O}}(\delta^{3}).$ (2.19b)
@@ 2.3.5 Governing Equations
To derive the ${\mathcal{O}}(\delta^{3})$ averaged equations, we follow a procedure similar to the one used for lower-order terms. Specifically, we obtain the ${\mathcal{O}}(\delta^{3})$ governing equations by combining $\delta$ times (2.14), $\delta^{2}$ times (2.19), and $\delta^{3}$ times the ${\mathcal{O}}(\delta^{3})$ equations (the detailed expressions are omitted here due to their complexity). We then simplify the expression and combine terms using (2.13). Afterward, we convert all time derivatives to spatial derivatives, except for $\rho_{t}$ from the first equation and $q_{t}$ from the second, to derive evolution equations analogous to those of isentropic gas equations. Utilizing the averaged equations to replace higher-order derivatives in time and the mixed derivatives. We then obtain:
@@
$\ \ -\frac{2C_{3}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\underline{q}\,\underline{\rho}\,\underline{\rho}_{x}+\left(\frac{2C_{8}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}-\frac{2C_{4}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right)\underline{q}_{x}\underline{\rho}_{x}+\left(-\frac{2C_{13}C_{1}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle^{2}}+\frac{2C_{8}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}-\frac{2C_{4}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right)\underline{q}\,\underline{\rho}_{xx}\Biggr{]}+{\mathcal{O}}(\delta^{4}),$
@@
$\ \ +\left(\frac{2C_{7}P^{\prime}}{\langle{a^{-1}}\rangle^{3}}+\frac{C_{1}P^{\prime\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}-\frac{2C_{1}C_{13}P^{\prime}}{\langle{a^{-1}}\rangle^{3}\langle{a}\rangle}\right)\underline{\rho}_{x}^{2}+\frac{C_{1}P^{\prime\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}\underline{\rho}\,\underline{\rho}_{xx}+\left(\frac{C_{11}P^{\prime}}{\langle{a^{-1}}\rangle^{3}\langle{a}\rangle}-\frac{C_{2}P^{\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}\right)\underline{\rho}_{xxx}\Biggr{]}+{\mathcal{O}}(\delta^{4}),$
where the coefficients are provided in Appendix A.
@@ 3 Linear Dispersion Relation
$\underline{\rho}_{t}$ $=\alpha_{1}\underline{q}_{x}+\alpha_{2}\underline{q}_{xx}+\alpha_{3}\underline{q}\,\underline{\rho}_{x}+\alpha_{4}\underline{q}^{2}\underline{q}_{x}+\alpha_{5}\underline{q}_{xxx}+\alpha_{6}\underline{q}\,\underline{\rho}\,\underline{\rho}_{x}+\alpha_{7}\underline{q}_{x}\,\underline{\rho}_{x}+\alpha_{8}\underline{q}\,\underline{\rho}_{xx},$ (3.2a)
$\underline{q}_{t}$ $=\beta_{1}\underline{\rho}_{x}+\beta_{2}\underline{q}\,\underline{q}_{x}+\beta_{3}\underline{\rho}\,\underline{\rho}_{x}+\beta_{4}\underline{\rho}_{xx}+\beta_{5}\underline{q}\,\underline{\rho}\,\underline{q}_{x}+\beta_{6}\underline{q}_{x}^{2}+\beta_{7}\underline{q}\,\underline{q}_{xx}+\beta_{8}\underline{q}^{2}\underline{\rho}_{x}$ (3.2b)
$\ \ +\beta_{9}\underline{\rho}_{x}^{2}+\beta_{10}\underline{\rho}\,\underline{\rho}_{xx}+\beta_{11}\underline{\rho}_{xxx}.$
@@
$\overline{\rho}_{t}$ $=\alpha_{1}\overline{q}_{x}+\alpha_{2}\overline{q}_{xx}+\alpha_{5}\overline{q}_{xxx},$ (3.3a)
$\overline{q}_{t}$ $=\beta_{1}\overline{\rho}_{x}+\beta_{3}\rho^{(0)}\overline{\rho}_{x}+\beta_{4}\overline{\rho}_{xx}+\beta_{10}\rho^{(0)}\overline{\rho}_{xx}+\beta_{11}\overline{\rho}_{xxx}.$ (3.3b)
@@
$-i\omega\overline{\rho}$ $=ik\alpha_{1}\overline{q}-k^{2}\alpha_{2}\overline{q}-ik^{3}\alpha_{5}\overline{q},$ (3.5a)
$-i\omega\overline{q}$ $=ik\beta_{1}\overline{\rho}+ik\beta_{3}\rho^{(0)}\overline{\rho}-k^{2}\beta_{4}\overline{\rho}-k^{2}\beta_{10}\rho^{(0)}\overline{\rho}-ik^{3}\beta_{11}\overline{\rho},$ (3.5b)
@@
$\begin{bmatrix}\omega&\hfill k\alpha_{1}+ik^{2}\alpha_{2}-k^{3}\alpha_{5}\\ k\beta_{1}+k\beta_{3}\rho^{(0)}+ik^{2}\beta_{4}+ik^{2}\beta_{10}\rho^{(0)}-k^{3}\beta_{11}\hfill&\omega\end{bmatrix}\begin{bmatrix}\overline{\rho}\\ \overline{q}\end{bmatrix}:=M\begin{bmatrix}\overline{\rho}\\ \overline{q}\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}.$
@@
$\omega(k)=\pm|k|\sqrt{\left(\alpha_{1}+ik\alpha_{2}-k^{2}\alpha_{5}\right)\left(\beta_{1}+\beta_{3}\rho^{(0)}+ik\beta_{4}+ik\beta_{10}\rho^{(0)}-k^{2}\beta_{11}\right)}.$ (3.6)
@@
$\underline{q}_{txx}$ $=-\frac{P^{\prime}(\underline{\rho}^{(0)})}{\langle{a^{-1}}\rangle}\underline{\rho}_{xxx}\iff\underline{\rho}_{xxx}=-\frac{\langle{a^{-1}}\rangle}{P^{\prime}(\underline{\rho}^{(0)})}\underline{q}_{txx}$ (3.7b)
@@
$\overline{\rho}_{t}$ $=\alpha_{1}\overline{q}_{x}+\alpha_{2}\overline{q}_{xx}\overline{q}_{x}+\alpha_{5b}\overline{\rho}_{xxt},$ (3.8a)
$\overline{q}_{t}$ $=\beta_{1}\overline{\rho}_{x}+\beta_{3}\rho^{(0)}\overline{\rho}_{x}+\beta_{4}\overline{\rho}_{xx}+\beta_{10}\rho^{(0)}\overline{\rho}_{xx}+\beta_{11b}\overline{q}_{xxt},$ (3.8b)
where
$\alpha_{5b}=\delta^{2}\left(\frac{-C_{9}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}+\frac{C_{2}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right),~{}~{}~{}~{}\beta_{11b}=\delta^{2}\left(\frac{-C_{11}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}+\frac{C_{2}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right)$
and the remaining coefficients are as given in the appendix. Therefore, the dispersion relation is
$\omega(k)=\pm|k|\sqrt{\frac{\alpha_{1}\left(\beta_{1}+\beta_{3}\rho^{(0)}\right)}{\left(1+\alpha_{5b}k^{2}\right)\left(1+\beta_{11b}k^{2}\right)}}$ (3.9)
The dispersion relation is stable when $\mathrm{Im}(\omega(k))=0$. Since the product of $\alpha_{1}$ and $\left(\beta_{1}+\beta_{3}\rho^{(0)}\right)$ is always positive for any $\rho^{(0)}\geq 0$, we can conclude that the condition for stability is satisfied. Specifically, $\alpha_{1}$, $\beta_{1}$, and $\beta_{3}$ depend only on $\langle{a}\rangle$ and the derivatives of the pressure, all of which are positive. For the denominator of the expression, $\alpha_{5b}$ and $\beta_{11b}$ are always positive for any periodic function, further ensuring stability. For the numerator, we can show that both $C_{9}$ and $C_{11}$ are negative. Starting with $C_{9}$, we have:

### v2

@@ 1 Introduction
The main goal of this work is to construct an effective medium that represents the essential characteristics of the original microstructured medium. To accomplish this, we develop a constant-coefficient (homogenized) approximation, which effectively describes the behavior of long-wavelength solutions for a system that describes the scenario of a gas propagating in a narrow pipe with a periodically-varying cross-sectional area. This approximation method draws upon the approach developed by Yong and co-authors [18, 8], and more recently was applied to the shallow water waves over periodically-varying bathymetry [17, 19], as well as the compressible Euler equations [20]. The constant-coefficient equations we derive have dispersive terms, demonstrating that waves in this context exhibit effective dispersion and allowing us to quantify the dispersive effects. While the formal process employed herein is very similar to that in the works just mentioned, the analysis here is more technical due to the combined presence of variable coefficients and a variable non-conservative term (see Equation (2.1)). The resulting homogenized equations are, correspondingly, more complicated than in any of the previous applications.
@@
Code to reproduce all simulation results in this paper can be found online [^1] https://github.com/laila-16/Homogenization_of_Isentropic_Gas_Equations.
@@ 2.2 Averaging Operators
  • $\langle{b(y)}\rangle=\int^{1}_{0}b(y)dy.$ This averaging operator gives the average value of a function over one period. The result of this operator is a function independent of $y$.
@@ 2.3.3 ${\mathcal{O}}(\delta^{1})$
$0$ $=a\underline{\rho}^{(1)}_{t}+\underline{q}^{(1)}_{x}+q^{(2)}_{y}$, (2.10a)
$0$ $=-\frac{a^{-2}a_{y}}{\underline{\rho}^{(0)}}(\underline{q}^{(1)})^{2}+\underline{q}^{(1)}_{t}+aP^{\prime}(\underline{\rho}^{(0)})\left(\underline{\rho}^{(1)}_{x}+\rho^{(2)}_{y}\right)$. (2.10b)
@@
To remove the $y$ dependence and avoid secular terms we apply the averaging operator $\langle{.}\rangle$. We obtain:
@@
The average of the term $\frac{a^{-3}a_{y}(\underline{q}^{(1)})^{2}}{\underline{\rho}^{(0)}P^{\prime}(\underline{\rho}^{(0)})}$ vanishes because $a(y)$ is periodic in $y$ and hence $\int_{\delta}a^{-3}a_{y}dy=0$. Notice that (2.12) is just the linearized and averaged form of the original system (2.2). We define the averaged variables
@@
$\underline{\rho}_{t}$ $=-\frac{\underline{q}_{x}}{\langle{a}\rangle}+{\mathcal{O}}(\delta),$ (2.14a)
$\underline{q}_{t}$ $=-\frac{P^{\prime}(\underline{\rho}^{(0)})}{\langle{a^{-1}}\rangle}\underline{\rho}_{x}+{\mathcal{O}}(\delta).$ (2.14b)
@@ 2.3.4 ${\mathcal{O}}(\delta^{2})$
$0$ $=a\rho^{(2)}_{t}+q^{(2)}_{x}+q^{(3)}_{y}$, (2.16a)
$0$ $=q^{(2)}_{t}+a^{-2}a_{y}\left(-\frac{2q^{(1)}q^{(2)}}{\rho^{(0)}}+\frac{(q^{(1)})^{2}\rho^{(1)}}{(\rho^{(0)})^{2}}\right)+2a^{-1}\left(\frac{q^{(1)}q^{(2)}_{y}+q^{(1)}q^{(1)}_{x}}{\rho^{(0)}}\right)$
@@
Applying the averaging operator $\langle{.}\rangle$ to the system above in order to remove the $y$ dependence yields the following averaged system:
@@
$\underline{\rho}_{t}$ $=-\frac{\underline{q}_{x}}{\langle{a}\rangle}+\delta\left(-\frac{2\langle{a\llbracket{a^{-3}a_{y}}\rrbracket}\rangle}{\langle{a^{-1}}\rangle\underline{\rho}^{(0)}\langle{a}\rangle}\underline{q}\,\underline{\rho}_{x}-\frac{\langle{a\llbracket{a^{-1}}\rrbracket}\rangle}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}\underline{q}_{xx}\right)+{\mathcal{O}}(\delta^{2}),$ (2.19a)
$\underline{q}_{t}$ $=-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}\underline{\rho}_{x}+\delta\Biggr[\left(\frac{-2}{\underline{\rho}^{(0)}\langle{a}\rangle}+\frac{2\langle{a^{-3}a_{y}\llbracket{a}\rrbracket}\rangle}{\langle{a^{-1}}\rangle\underline{\rho}^{(0)}\langle{a}\rangle}\right)\underline{q}\,\underline{q}_{x}-\frac{P^{\prime\prime}}{\langle{a^{-1}}\rangle}\underline{\rho}\,\underline{\rho}_{x}+\frac{\langle{a^{-1}\llbracket{a}\rrbracket}\rangle P^{\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}\underline{\rho}_{xx}\Biggl]+{\mathcal{O}}(\delta^{2}).$ (2.19b)
@@ 2.3.5 Governing Equations
To derive the ${\mathcal{O}}(\delta^{3})$ averaged equations, we follow a procedure similar to the one used for lower-order terms. Specifically, we obtain the ${\mathcal{O}}(\delta^{3})$ governing equations by combining $\delta$ times (2.14), $\delta^{2}$ times (2.19), and $\delta^{3}$ times the ${\mathcal{O}}(\delta^{3})$ equations (the detailed expressions are omitted here due to their complexity). We then simplify the expression and combine terms using (2.13).
The resulting equations contain terms with high-order time derivatives. A linear dispersion analysis of the equations reveals that they are unstable. This is a known issue with certain homogenization techniques like that employed herein; see [7] for a general discussion and [8, 17] for similar examples. What is recommended in [7], and also employed in [8], is to exchange all high-order time derivatives (i.e., all except those appearing in the linear evolution term for each equation) for space derivatives, by using equality of mixed partial derivatives and keeping the same formal order of accuracy in $\delta$. This approach leads to a system that is also linearly unstable, but in a weaker sense – it exhibits instability only for high-wavenumber modes, and consequently it is possible to use it for numerical simulations (as is done in [8]) as long as the mesh is not too fine.
This yields
@@
$\ \ -\frac{2C_{3}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\underline{q}\,\underline{\rho}\,\underline{\rho}_{x}+\left(\frac{2C_{8}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}-\frac{2C_{4}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right)\underline{q}_{x}\underline{\rho}_{x}+\left(-\frac{2C_{13}C_{1}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle^{2}}+\frac{2C_{8}}{\langle{a^{-1}}\rangle\langle{a}\rangle^{2}}-\frac{2C_{4}}{\langle{a^{-1}}\rangle\langle{a}\rangle}\right)\underline{q}\,\underline{\rho}_{xx}\Biggr]+{\mathcal{O}}(\delta^{3}),$
@@
$\ \ +\left(\frac{2C_{7}P^{\prime}}{\langle{a^{-1}}\rangle^{3}}+\frac{C_{1}P^{\prime\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}-\frac{2C_{1}C_{13}P^{\prime}}{\langle{a^{-1}}\rangle^{3}\langle{a}\rangle}\right)\underline{\rho}_{x}^{2}+\frac{C_{1}P^{\prime\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}\underline{\rho}\,\underline{\rho}_{xx}+\left(\frac{C_{11}P^{\prime}}{\langle{a^{-1}}\rangle^{3}\langle{a}\rangle}-\frac{C_{2}P^{\prime}}{\langle{a^{-1}}\rangle^{2}\langle{a}\rangle}\right)\underline{\rho}_{xxx}\Biggr]+{\mathcal{O}}(\delta^{3}),$
where the coefficients are provided in Appendix A. We remark that for a wide class of periodic functions $a(x)$ (specifically, those that are translation-even in the terminology of [17, Appendix A]) we have $C_{1}=C_{7}=0$, so that several of the coefficients above vanish. In the examples in Section 4 we also choose such functions $a(x)$, and in some of the following analysis we will consider this special, slightly simpler case.
Similar to the other applications mentioned above, these equations are unstable for sufficiently high wavenumbers (see our analysis below). In order to obtain a fully stable system of equations, we follow what is done in [17] and was suggested by observations dating back to Whitham [22, Section 13.11]. Namely, we exchange some time derivatives for space derivatives by using equality of mixed partial derivatives, but we keep exactly one $t$-derivative in each higher-order linear term, thus obtaining a system that is linearly stable for all wavenumbers. The resulting system can be written (introducing symbols for the lengthy coefficients, which are also defined Appendix A):
$\underline{\rho}_{t}$ $=-\frac{1}{\langle{a}\rangle}\underline{q}_{x}+\delta\alpha_{3}\underline{q}\,\underline{\rho}_{x}+\delta^{2}\left(\alpha_{4}\underline{q}^{2}\underline{q}_{x}-\langle{a}\rangle\alpha_{5}\underline{\rho}_{xxt}+\alpha_{6}\underline{q}\,\underline{\rho}\,\underline{\rho}_{x}+\alpha_{7}\underline{q}_{x}\underline{\rho}_{x}+\alpha_{8}\underline{q}\,\underline{\rho}_{xx}\right),$ (2.22a)
$\underline{q}_{t}$ $=-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}\underline{\rho}_{x}+\delta\left(\beta_{2}\underline{q}\,\underline{q}_{x}+\beta_{3}\underline{\rho}\,\underline{\rho}_{x}\right)+\delta^{2}\Bigl(\beta_{5}\underline{q}\,\underline{\rho}\,\underline{q}_{x}+\beta_{6}\underline{q}_{x}^{2}+\beta_{7}\underline{q}\,\underline{q}_{xx}+\beta_{8}\underline{q}^{2}\underline{\rho}_{x}$ (2.22b)
$\ \ +\beta_{9}\underline{\rho}_{x}^{2}-\frac{\langle{a^{-1}}\rangle}{P^{\prime}}\beta_{11}\underline{q}_{xxt}\Bigr).$
We conclude this session with a remark regarding the ansatz (2.5). It is possible to instead suppose a priori that the waves of interest are quadratically small with respect to $\delta$. This leads to a similar set of equations except that the power of $\delta$ multiplying each term is different. In particular, with such an ansatz the quadratic nonlinear terms and the third-derivative terms appear at the same order as one another.
@@ 3 Linear Dispersion Relation
$\underline{\rho}_{t}$ $=-\frac{1}{\langle{a}\rangle}\underline{q}_{x}+\delta\alpha_{3}\underline{q}\,\underline{\rho}_{x}+\delta^{2}\left(\alpha_{4}\underline{q}^{2}\underline{q}_{x}+\alpha_{5}\underline{q}_{xxx}+\alpha_{6}\underline{q}\,\underline{\rho}\,\underline{\rho}_{x}+\alpha_{7}\underline{q}_{x}\,\underline{\rho}_{x}+\alpha_{8}\underline{q}\,\underline{\rho}_{xx}\right),$ (3.2a)
$\underline{q}_{t}$ $=-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}\underline{\rho}_{x}+\delta\left(\beta_{2}\underline{q}\,\underline{q}_{x}+\beta_{3}\underline{\rho}\,\underline{\rho}_{x}\right)+\delta^{2}\Bigl(\beta_{5}\underline{q}\,\underline{\rho}\,\underline{q}_{x}+\beta_{6}\underline{q}_{x}^{2}+\beta_{7}\underline{q}\,\underline{q}_{xx}+\beta_{8}\underline{q}^{2}\underline{\rho}_{x}+\beta_{9}\underline{\rho}_{x}^{2}$ (3.2b)
$\ \ +\beta_{11}\underline{\rho}_{xxx}\Bigr).$
@@
$\overline{\rho}_{t}$ $=-\frac{1}{\langle{a}\rangle}\overline{q}_{x}+\delta^{2}\alpha_{5}\overline{q}_{xxx},$ (3.3a)
$\overline{q}_{t}$ $=-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}\overline{\rho}_{x}+\delta\beta_{3}\rho^{(0)}\overline{\rho}_{x}+\delta^{2}\beta_{11}\overline{\rho}_{xxx}.$ (3.3b)
@@
$-i\omega\overline{\rho}$ $=-\frac{1}{\langle{a}\rangle}ik\overline{q}-ik^{3}\delta^{2}\alpha_{5}\overline{q},$ (3.5a)
$-i\omega\overline{q}$ $=-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}ik\overline{\rho}+ik\delta\beta_{3}\rho^{(0)}\overline{\rho}-ik^{3}\delta^{2}\beta_{11}\overline{\rho},$ (3.5b)
@@
$\begin{bmatrix}\omega&\hfill-\frac{1}{\langle{a}\rangle}k-k^{3}\delta^{2}\alpha_{5}\\ -\frac{P^{\prime}}{\langle{a^{-1}}\rangle}k+k\delta\beta_{3}\rho^{(0)}-k^{3}\delta^{2}\beta_{11}\hfill&\omega\end{bmatrix}\begin{bmatrix}\overline{\rho}\\ \overline{q}\end{bmatrix}:=M\begin{bmatrix}\overline{\rho}\\ \overline{q}\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}$.
@@
$\omega(k)=\pm|k|\sqrt{\left(-\frac{1}{\langle{a}\rangle}-k^{2}\delta^{2}\alpha_{5}\right)\left(-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}+\delta\beta_{3}\rho^{(0)}-k^{2}\delta^{2}\beta_{11}\right)}.$ (3.6)
@@
$\underline{q}_{txx}$ $=-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}\underline{\rho}_{xxx}\iff\underline{\rho}_{xxx}=-\frac{\langle{a^{-1}}\rangle}{P^{\prime}}\underline{q}_{txx}$ (3.7b)
@@
$\overline{\rho}_{t}$ $=-\frac{1}{\langle{a}\rangle}\overline{q}_{x}-\langle{a}\rangle\delta^{2}\alpha_{5}\overline{\rho}_{xxt},$ (3.8a)
$\overline{q}_{t}$ $=-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}\overline{\rho}_{x}+\delta\beta_{3}\rho^{(0)}\overline{\rho}_{x}-\frac{\langle{a^{-1}}\rangle}{P^{\prime}}\delta^{2}\beta_{11}\overline{q}_{xxt}.$ (3.8b)
Therefore, the dispersion relation is
$\omega(k)=\pm|k|\sqrt{\frac{-\frac{1}{\langle{a}\rangle}\left(-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}+\delta\beta_{3}\rho^{(0)}\right)}{\left(1-\langle{a}\rangle\delta^{2}\alpha_{5}k^{2}\right)\left(1-\frac{\langle{a^{-1}}\rangle}{P^{\prime}}\delta^{2}\beta_{11}k^{2}\right)}}$ (3.9)
The dispersion relation is stable when $\mathrm{Im}(\omega(k))=0$. Since the product of $\frac{1}{\langle{a}\rangle}$ and $\left(-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}+\delta\beta_{3}\rho^{(0)}\right)$ is always positive for any $\rho^{(0)}\geq 0$, we can conclude that the condition for stability is satisfied. Specifically, $\frac{1}{\langle{a}\rangle}$, $\frac{P^{\prime}}{\langle{a^{-1}}\rangle}$, and $\beta_{3}$ depend only on $\langle{a}\rangle$ and the derivatives of the pressure, all of which are positive. For the denominator of the expression, $\langle{a}\rangle\alpha_{5}$ and $\frac{\langle{a^{-1}}\rangle}{P^{\prime}}\beta_{11}$ are always positive for any periodic function, further ensuring stability. For the numerator, we can show that both $C_{9}$ and $C_{11}$ are negative. Starting with $C_{9}$, we have:
@@
In Figure 2, we compare the dispersion relations for the specific cross-sections under consideration. We observe that at large wave numbers, the dispersion relation given by (3.6) grows significantly, which can impact the stability of the numerical solution. To analyze the agreement between the two models, we perform a Taylor expansion of both dispersion relations.
For both dispersion relations (3.6) and (3.9), the Taylor expansion gives
$\omega(k)=\sqrt{-\frac{1}{\langle{a}\rangle}(-\frac{P^{\prime}}{\langle{a^{-1}}\rangle}+\delta\beta_{3}\rho_{0})k^{2}}+\mathcal{O}(k^{3}),$
since the $\mathcal{O}(k^{2})$ terms vanish in each case. Thus, the two dispersion relations coincide up to the included order, and their difference satisfies $\omega_{\text{(3.6)}}(k)-\omega_{\text{(3.9)}}(k)=\mathcal{O}(k^{3})$.
Figure 2: Dispersion relations with piecewise-constant (left) and sinusoidal (right) cross-sections. Solid lines correspond to KdV-type dispersion (3.6), and dashed lines to BBM-type dispersion (3.9). Note that for (3.6), we truncate the plot at the point where $\omega(k)$ becomes complex.